OpenStudy (kate_kat):

PLZ HELP! Test each set of equivalent ratios to find which set of values describes the proportional relationship of the two quantities. 17 ounces of nuts to 22.1 ounces of raisins A) (5, 7)(60, 84)(47, 65.8) B) (36, 48.6)(2, 2.7)(12, 16.2) C) (3, 3.9)(65, 84.5)(24, 31.2) D) (11, 13.2)(99, 118.8)(20, 24)

OpenStudy (kate_kat):

@zepdrix can you help?

OpenStudy (welshfella):

First divide 22.1 by 17 What do you get.p

OpenStudy (kate_kat):

1.3

OpenStudy (welshfella):

Right Now check if you get 1.3 when you divide numbers in the given choices

OpenStudy (kate_kat):

so i would divid 5 and 7 then 60 and 84 ect.

OpenStudy (kate_kat):

is that right?

OpenStudy (justin234):

Nice job explaining @welshfella

OpenStudy (welshfella):

Yea 7/5 = .?

OpenStudy (kate_kat):

1.4

OpenStudy (welshfella):

ty

OpenStudy (justin234):

*appluase*

OpenStudy (welshfella):

Ok so it can't be A. right?

OpenStudy (kate_kat):

it wouldn't be any of them none of then equal 1.3

OpenStudy (kate_kat):

sooo, what do i do if none of them are right???

OpenStudy (justin234):

one of them matches

OpenStudy (kate_kat):

i just divided all of them up and none = 1.3 tho

OpenStudy (justin234):

check ur original answers and see if they match the answers here

OpenStudy (kate_kat):

so i check to see if they equal 1.3?

OpenStudy (justin234):

yea

OpenStudy (kate_kat):

ok hold on

OpenStudy (justin234):

Just in case ur answers written here are written wrong

OpenStudy (kate_kat):

ok i did it and none of them equal 1.3 :(

OpenStudy (kate_kat):

what is the answer u got because i literally got nothing

OpenStudy (justin234):

ok then it must be how u and welshfella problem solving

OpenStudy (kate_kat):

@welshfella what was the answer u got?

OpenStudy (justin234):

try ur second idea 1.4

OpenStudy (justin234):

on the original answers

OpenStudy (kate_kat):

1.4 was A it wouldnt be the answer it has to equal 1.3

OpenStudy (kate_kat):

cuz 17/22.1=1.3

OpenStudy (justin234):

right so it has to 1.3 but it doesnt work ur sayin?

OpenStudy (kate_kat):

yes im saying that none of the answers equal 1.3 ...

OpenStudy (kate_kat):

i think im just gonna guess and tell you what the answer was cuz im just comfused...

OpenStudy (justin234):

I am not confused too idk what @welshfella thinks but im confused

OpenStudy (kate_kat):

i dont know how but it was C

OpenStudy (kate_kat):

By creating a table of values you can check to see whether each set of values has a multiplicative relationship to the next set. The ratio of nuts to raisins is 1 to 1.3.

OpenStudy (welshfella):

Try option C

OpenStudy (kate_kat):

yeah but how did u get that welshfella, cuz when i divided i didnt get anything

OpenStudy (welshfella):

3.9/3=?

OpenStudy (kate_kat):

i thought it was 3/3.9

OpenStudy (kate_kat):

thx for your help @welshfella

OpenStudy (kate_kat):

thank you too @Justin234